OpenStudy (anonymous):

solve: 5/y-4-8/y-4=2/y2-16

OpenStudy (anonymous):

\[\frac{5}{-4+y}-\frac{8}{-4+y}\text{=}\frac{2}{-16+y^2} \]Move the RHS to the LHS changing the sign.\[\frac{5}{-4+y}-\frac{8}{-4+y}-\frac{2}{-16+y^2}=0 \]Combine the fractions and simplify.\[\frac{-14-3 y}{(-4+y) (4+y)}= 0\]\[-14-3 y=0\]\[y=-\frac{14}{3} \]